OpenStudy (anonymous):

if you have 15 questions and you miss 9 what is you grade...:]] thanks to all who answer

OpenStudy (phi):

if you miss 9 out of 15, how many did you get correct?

OpenStudy (anonymous):

6..so that means a 60

OpenStudy (phi):

if you get 6 out of 15 correct , you write that as \[\frac{6}{15} \] you can simplify by dividing top and bottom by 3 to get \[\frac{2}{5} \] you can change this to a fraction by dividing 5 into 2 to get 0.4 you can change the decimal to a percent by multiplying it by 100 you get 0.4*100= 40%